OpenStudy (anonymous):

Determine whether the ordered pair is a solution to the given linear inequality: x-7y≥ -3 (a) (4,1) , (b) (9,2), (c) (0,0) , (d) (-16,-2)

OpenStudy (azureilai):

You can plug everything in and determine which pair is true.

OpenStudy (anonymous):

Can you solved it

OpenStudy (azureilai):

4-7(1)=-3 9-7(2)=-5 0-7(0)=0 -16-7(-2)=2 It would be a

OpenStudy (azureilai):

medal please

OpenStudy (anonymous):

you didnt finish it

OpenStudy (anonymous):

 f4g=m+x-c